OpenStudy (anonymous):

Help please! The root x = 1 has multiplicity 3 for the function f(x)=x^3-x^2-x+1. True or false

OpenStudy (campbell_st):

well if x = 1 is a root then (x -1) is a factor multiplicity means that its (x -1)(x -1)(x -1) = (x -1)^3 now distribute the 1st pair of (x -1)s and you get (x^2 -2x + 1)(x -1) I'll let you finish and check... hope it helps...

OpenStudy (anonymous):

Based on that, I think it's false! Is that correct? @campbell_st

OpenStudy (ybarrap):

Using @campbell_st 's fine description then multiply out, we get: (x-1)^3=x^3-3 x^2+3 x-1. Compare this to your given equation, if they are equal then answer is True.

OpenStudy (campbell_st):

well using a little savvy... the costant term when multiplied out will be -1 and your function has a constant of +1 so it would be false..